Nominal gross domestic product for the United States was $7, 661.6 billion in 1996 and $8,110.9 billion in 1997. The GDP deflator was 109.5 in 1996 and 111.6 in 1997. Given the above data, complete the following questions. SHow the work.

  1. What was the 1997 real GDP expressed in 1996 prices?
  2. What were the 1996 and 1997 GDPs expressed in base year prices?
  3. What was the percent change in the nominal GDP between 1996 and 1997?
  4. What was the percent change in real GDP between 1996 and 1997?
  5. What was inflation between 1996 and 1997?
1
Expert's answer
2020-11-10T10:38:47-0500

1) "\bold {Answer}"

1997 real GDP in 1996 prices "= \$7,958.28 \space bn"


"\bold {Solution}"


1997 real GDP in 1996 prices

"= \dfrac {1997 \space Nominal \space GDP}{1997 \space GDP \space deflator}× 1996 \space GDP \space deflator"


"= \dfrac {\$8,110.9 \space bn}{111.6}×109.5"


"= \$7,958.2755376 \space bn"


"= \bold {\$7,958.28 \space bn}"



2) "\bold {Answers}"


1996 real GDP "= \$6,996.89 \space bn"


1997 real GDP "= \$7,267.83 \space bn"


"\bold {Solutions}"


GDP in base year prices

= Real GDP

"= \dfrac {Nominal \space GDP}{GDP \space Deflator }"



"1996 = \dfrac {\$7,761.6 \space bn}{1.095}"


"= \$6,996.8949772 \space bn"


"= \bold {\$6,996.89 \space bn}"



"1997 = \dfrac {\$8110.9 \space bn}{1.116}"


"= \$7,267.8315412 \space bn"


"= \bold {\$7,267.83 \space bn}"




3) "\bold {Answer}"


Percentage change in nominal GDP

"= 5.86\%"


"\bold {Solution}"


Percentage change in nominal GDP

"= \dfrac {1997 \space nominal \space GDP - 1996 \space nominal \space GDP}{1996 \space nominal \space GDP} ×100\%"


"= \dfrac {\$8,110.9 \space bn - \$7,661.6 \space bn}{\$7,661.6}×100\%"


"= \dfrac {\$449.3 \space bn}{\$7,661.6 \space}×100\%"


"= 5.86431032\%"


"= \bold {5.86\%}"




4) "\bold {Answer}"


Percentage change in real GDP

"= 3.86\%"


"\bold {Solution}"


Percentage change in real GDP

"= \dfrac {1997 \space real \space GDP -1996 \space real \space GDP}{1996 \space real \space GDP} × 100\%"

"= \dfrac {\$7,267.83 \space bn-\$6,996.89 \space bn}{\$6,996.89}×100\%"


"= \dfrac {\$270.94}{\$6,996.89}×100\%"


"= 3.86457411\%"


"= \bold {3.86\%}"




5) "\bold {Answer}"


Inflation rate "= 1.92\%"



"\bold {Solution}"


Inflation rate

= percentage change in GDP deflator

"=\dfrac {1997 \space deflator - 1996 \space deflator } {1996 \space deflator }×100\%"

"= \dfrac {111.6-109.5}{109.5}×100\%"


"= \dfrac {2.1}{108.5}×100\%"


"= 1.91780821\%"


"= \bold {1.92\%}"